Circular No. 8499

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005av IN NGC 6943 Further to IAUC 8496, L. A. G. Monard reports his discovery of an

apparent supernova (mag ∼ 15.4 ± 0.1) on unfiltered CCD images taken on Mar. 24.077 and 24.983 UT with a 0.30-m reflector near Pretoria, South Africa. The new object is located in a spiral arm of NGC 6943 at α = 20^{h}44^{m}37^{s}.58, δ = −68^{o}45′10′′.6 (equinox 2000.0), which is 22′′ east and 18′′ south of the nucleus of the galaxy. Nothing is visible at this location on Monard’s image taken on Mar. 5.133 (limiting mag 17.5) or on an image from the Digitized Sky Survey (limiting red mag 20.5). SN 1998ew also occurred in this galaxy.

SUPERNOVA 2005aw IN IC 4837A Further to IAUC 8488, C. Jacques and E. Pimentel report their dis-

covery of an apparent supernova (mag 15.3 ± 0.2) on unfiltered CCD images taken on Mar. 24.27 and 25.25 UT with a 0.30-m Schmidt-Cassegrain reflector at Belo Horizonte, Brazil. The new object is located at α = 19^{h}15^{m}17^{s}.44, δ = −54^{o}08′24′′.9 (equinox 2000.0), which is 13′′ east and 30′′ south of the center of IC 4837A. Nothing is visible at this location on a CCD image taken on Mar. 7.32 (limiting mag 17.9) or on a 1976 Palomar Sky Survey plate (limiting red mag 20.0).

SUPERNOVA 2005ax IN NGC 1544 Further to IAUC 8492, T. Puckett and D. George report the discovery

of an apparent supernova (mag 14.5) on an unfiltered CCD frame taken with the 0.60-m automated supernova patrol telescope on Mar. 26.12 UT. The new object was confirmed on multiple combined CCD frames taken by D. George (Ottawa, Canada) with a 0.35-m reflector on Mar. 27.31 (at mag 14.5). The new object is located at α = 5^{h}02^{m}38^{s}.11, δ = +86^{o}13′22′′.2 (equinox 2000.0), which is 2′′ east and 2′′ north of the center of NGC 1544. Nothing is visible at this location on images taken by Puckett on 2002 Apr. 5 and 2003 Feb. 11 (limiting mag ∼ 20.0).

COMET C/1996 V2 (SOHO) Continuation to IAUC 8498 (another Marsden-group member):

Comet 1996 UT α_{2000} δ_{2000} Inst. F MPEC
C/1996 V2 Nov. 11.986 14^{h}59.^{m}3 −15^{o}54′ C3 RK 2005-F31
2005 March 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green